OpenStudy (revircs):

Consider the function f(x)=cosh(3x). For the following questions, express your answers in terms of exponential functions. a) Find the local linearization of f(x) near x=1 b) Does the approximation overestimate or underestimate f(2)? I really have no idea where to start with this.

OpenStudy (16shuston):

How I would start this and Im not 100% sure but i would plug 1 into x

OpenStudy (16shuston):

It should look like cos(3*1) then plug that into your calculator for part A For part B take and do the same thing except with the number 2

OpenStudy (baru):

this is how u start: linear approximation \[\Delta y= f'(x_0) \Delta x\]

OpenStudy (revircs):

So, would I need to differentiate cosh(x) first?

OpenStudy (baru):

yep

OpenStudy (revircs):

So that would come to \[\frac{ e^x-e^-x }{ 2 }\] ?

OpenStudy (baru):

i'm sorry xD i dont remember derivates of cosh or anything about hyperbolic functions, but i know what the question is asking for, i can help you with the steps

OpenStudy (revircs):

okay i know I got my derivative wrong, it is supposed to be a + instead of -, but what would I do from here?

OpenStudy (baru):

you evaluate f '(x) at x_0, here it says linearize near x=1, so x_0=1 substitute, get a number for f '(x_0)

OpenStudy (baru):

@Revircs following?

OpenStudy (revircs):

yeah, im just writing everything down for notes, sorry

OpenStudy (baru):

no prob :)

OpenStudy (revircs):

so after evaluating at x=1, I would get to \[\frac{ e^1+e^-1 }{ 2 }\]

OpenStudy (revircs):

or \[\frac{ e+1/e }{ 2 }\]

OpenStudy (baru):

ok, you can leave it that way for now, but later use a calculator, and actually find a number ok so now were done with f ' ( ) the question asks you to find f (2) using the linear approx, \(\Delta x\) represents change in x, so how much has x changed when we move from f(1) to f(2) ?

OpenStudy (revircs):

oh, okay. my professors usually have us leave it in exact form. So from there I would evaluate at f(x)=2 and find the difference?

OpenStudy (baru):

noo... you should realize that \(\Delta x=1\) substitute, f '(1) and \(\Delta x =1\) to find \(\Delta y \)

OpenStudy (revircs):

oh, so since the change in x is just 1, the change in y would be the same as the derivative in this case?

OpenStudy (baru):

yes

OpenStudy (baru):

i hope you are catching on to what linear approximation means, essentially, we are assuming that the graph of f(x) between the points x=1 and x=2 is a straight line with the slope f ' (1) 

OpenStudy (baru):

* sorry the two points marked on the x axis are 1 and 2

OpenStudy (baru):